# On $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ and Frege Systems

### v3

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). We further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. As a by-product, we reach the important result [Lin21] that $\mathcal{P}\neq\mathcal{NP}$ once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there are language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences in the area of proof complexity implied by our main outcome. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.
@@ 7.1 Lower Bounds for Frege Systems
Since nondeterministic Turing machines can simulate Frege proof (see [Bus99]) and by Theorem 7.1 any two Frege systems $p$-simulate each others, we can present the proof of Theorem 6 as follows:
@@ 8 Conclusions
One of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$ is basically the same as the one in the author’s [Lin21]. In other words, we first enumerate all polynomial-time ${\rm co}\mathcal{NP}$-machines, then by lazy-diagonalization we show that there exists a language $L_{d}$ accepted by a universal nondeterministic Turing machine running within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$, but by no ${\rm co}\mathcal{NP}$-machines. The next is to use essentially the same technique developed in author’s [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine is in fact in $\mathcal{NP}$, thus establishing the main theorem of this paper.
@@
As mentioned in the introduction Section, the question of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomial bounded, thus answering an open problem in [Pud08]. As pointed out in Remark 7.2, we did the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 6 is that nondeterministic Turing machines can simulate Frege proofs. But, we do not know whether the nondeterministic Turing machines can simulate Extended Frege proofs? For more detailed information about Extended Frege Systems, consult the reference [CR79, CN10, Kra95].

### v4

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there are language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which is in the area of proof complexity implied by our main outcome. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.
@@ 7.1 Lower Bounds for Frege Systems
Since nondeterministic Turing machines can simulate Frege proof (see e.g. [Bus99]) and by Theorem 7.1 any two Frege systems $p$-simulate each others, we can present the proof of Theorem 6 as follows:
@@ 8 Conclusions
One of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$ is basically the same as the one in the author’s recent work [Lin21]. In other words, we first enumerate all polynomial-time ${\rm co}\mathcal{NP}$-machines, then by lazy-diagonalization we show that there exists a language $L_{d}$ accepted by a universal nondeterministic Turing machine running within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$, but by no ${\rm co}\mathcal{NP}$-machines. The next is to use essentially the same technique developed in author’s work [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine is in fact in $\mathcal{NP}$, thus establishing the main theorem of this paper.
@@
As mentioned in the introduction Section, the question of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomial bounded, thus answering an open problem in [Pud08]. As pointed out in Remark 7.2, we did the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 6 is that nondeterministic Turing machines can simulate Frege proofs. But, we do not know whether the nondeterministic Turing machines can simulate Extended Frege proofs. For more detailed information about Extended Frege Systems, consult the reference [CR79, CN10, Kra95].